### v1

# High-energy Emission from Turbulent Electron-ion Coronae of Accreting Black Holes

Abstract
We develop a model of particle energization and emission from strongly turbulent black-hole coronae. Our local model is based on a set of 2D radiative particle-in-cell simulations with an electron-ion plasma composition, injection and diffusive escape of photons and charged particles, and self-consistent Compton scattering. We show that a radiatively compact turbulent corona generates extended nonthermal ion distributions, while producing X-ray spectra consistent with observations. As an example, we demonstrate excellent agreement with observed X-ray spectra of NGC 4151. The predicted emission spectra feature an MeV tail, which can be studied with future MeV-band instruments. The MeV tail is shaped by nonthermal electrons accelerated at turbulent current sheets. We also find that the corona regulates itself into a two-temperature state, with ions much hotter than electrons. The ions carry away roughly 60% to 70% of the dissipated power, and their energization is driven by a combination of shocks and reconnecting current sheets, embedded into the turbulent flow.
[…]
software: Tristan-MP v2 (Hakobyan et al., 2023)
show]daniel.groselj@kuleuven.be
I Introduction
Supermassive accreting black holes in active galactic nuclei (AGNs) are among the most powerful sources of hard X-rays (Laha et al., 2025; Kara & García, 2025). Based on their prevalence and physical properties, they are believed to be also among the prime sources of high-energy cosmic rays and neutrinos (Inoue et al., 2019; Murase et al., 2020; Padovani et al., 2024; Abbasi et al., 2025). The hard X-ray emission from AGNs is commonly attributed to a spatially compact region near the black hole, referred to as the “corona.” The dimensionless ratio of the coronal luminosity $L$ to its size $R$ defines the radiative compactness $\ell=4\pi(m_{\rm p}/m_{\rm e})(L/L_{\rm E})(R/R_{\rm g})^{-1}$ (Guilbert et al., 1983), where $L_{\rm E}$ is the Eddington limit, $R_{\rm g}$ is the gravitational radius, and $m_{\rm p}/m_{\rm e}$ is the proton-electron mass ratio. Many observed accreting black holes have $\ell\gg 1$, which implies that black-hole coronae are radiatively dense environments where matter and radiation are strongly coupled. Most notably, the high-compactness regime is characterized by moderate optical depths, fast radiative cooling, and (in the most extreme cases) electron-positron pair creation (e.g., Svensson, 1984; Fabian et al., 2015; Beloborodov, 2017).
A well-established view is that the coronal X-ray emission stems from the Compton scattering of low-energy seed photons off hot electrons with an effective temperature of about $100$ keV (Shapiro et al., 1976; Haardt & Maraschi, 1993). The rapid cooling experienced by the Comptonizing electrons points to an efficient form of coronal energy dissipation, which is needed to maintain the electrons at the observationally inferred temperatures.
Much insight into the physics of black-hole coronae can be gained by considering how the electrons are energized. The physics of electron energization involves collective plasma interactions, such as turbulence and/or magnetic reconnection (Petrosian, 2012; Lazarian et al., 2012; Guo et al., 2024; Sironi et al., 2025), which release the energy stored in the coronal magnetic fields and bulk motions into heat, nonthermal particles, and radiation. Thus, studies of plasma dissipation under the conditions that reproduce the observed coronal X-ray signal can place important constraints on the physics of the emission mechanism.
The same process which powers the electron energization, and in turn the X-ray emission, determines also the energetics of the coronal ions (i.e., protons). The composition of black-hole coronae is presently not well known; it could be either leptonic and dominated by pairs or of the hadronic type with a substantial amount of protons. In a number of sources, the compactness $\ell$ may not be high enough to sustain a pair-plasma-dominated state (Hinkle & Mushotzky, 2021). Models focusing on electron-ion compositions are particularly timely, given the recent neutrino observations by IceCube (Abbasi et al., 2025), which suggest that AGN coronae may be efficient proton accelerators. How the released energy is partitioned between the escaping photons and cosmic-ray protons can be investigated in particle-in-cell (PIC) simulations with explicit radiative transfer, which track the self-consistent interactions among photons, charged particles and their electromagnetic fields (Grošelj et al., 2024; Nättilä, 2024). However, existing studies focus mainly on pair plasma compositions.
In this paper, we study the partitioning of energy among photons and charged particles in strongly turbulent electron-ion coronae of accreting black holes. To this end, we perform the first PIC simulations with self-consistent Compton scattering in an electron-ion turbulent plasma. Alongside the numerical results we develop theoretical estimates. The predicted X-ray spectra are compared with observations of NGC 4151—one of the X-ray brightest AGNs (Ulrich, 2000) and a candidate neutrino source (Inoue & Khangulyan, 2023; Neronov et al., 2024; Abbasi et al., 2025; Murase et al., 2024). As we demonstrate below, the regime of electron-ion turbulence applicable to black-hole coronae features trans-sonic and trans-Alfvénic motions, leading to sporadic formation of magnetized shocks. The shocks, alongside reconnecting current sheets, act as drivers of particle heating and acceleration. The electron energization is largely offset by the rapid radiative cooling, which results in the formation of a two-temperature corona with ions much hotter than electrons ($T_{\rm i}\gg T_{\rm e}$). When the corona is sufficiently magnetized, we find that ions form extended nonthermal energy spectra. Our results have direct implications for the production of X-ray photons, cosmic rays and neutrinos in coronae of supermassive black holes.
II Theoretical Expectations
We consider a scenario where the coronal dissipation is mediated by magnetized turbulence in a weakly collisional electron-proton plasma. [^1] Our analytical estimates focus on the case where the plasma ions are protons with mass $m_{\rm i}\approx 1836\,m_{\rm e}$. However, the simulations (Sec. IV) employ an artificially reduced ion mass of $m_{\rm i}=144\,m_{\rm e}$ for computational convenience. Turbulently energized charged particles and photons escape the source via spatial diffusion, while at the same time low-energy thermal plasma and seed photons are resupplied into the domain. Energy released into the electrons is quickly passed onto photons and radiated away, while the fraction of energy received by the protons goes into heating and nonthermal acceleration. Below, we provide analytical estimates in support of this picture.
II.1 Regime of Turbulence
The global shape and multiphase structure of black-hole coronae, including the field line geometry, is presently the subject of ongoing investigations (e.g., Jiang et al., 2019; Chashkina et al., 2021; Liska et al., 2022, 2023; Scepi et al., 2024; Sridhar et al., 2025; Krawczynski & Hu, 2025). On the other hand, from the point of view of a local model, important constraints on the regime of turbulence can be obtained based on two aspects: i) the observationally inferred large radiative compactness $\ell\gg 1$, and ii) the presumed efficient particle acceleration. Alongside recent IceCube observations (Abbasi et al., 2025), hinting at proton acceleration, measurements of an extended MeV emission tail in the stellar-mass binary Cygnus X-1 (McConnell et al., 2002; Zdziarski et al., 2017) imply that black-hole coronae can also accelerate electrons into nonthermal power laws.
Amplitude of turbulent fluctuations
Recent studies of kinetic turbulence in local boxes with a mean magnetic field $B_{0}$ showed that efficient nonthermal particle acceleration strongly favors the large-amplitude turbulence regime with typical magnetic fluctuation strength $\delta B/B_{0}\gtrsim 1$ (Comisso & Sironi, 2018; Nättilä & Beloborodov, 2022; Vega et al., 2024). Since we aim to uncover regimes of efficient particle acceleration, we focus here on the case $\delta B/B_{0}\sim 1$.
Radiative compactness
For our model, the compactness $\ell$ can be expressed as follows. In steady state, the power channeled into escaping radiation per unit volume of a turbulent electron-ion plasma is $\dot{U}_{\rm esc}\simeq(1-q_{\rm i})B_{0}^{2}v_{\rm A}/(8\pi s_{0})$, where $q_{\rm i}$ is the ion heating fraction, $s_{0}$ is the turbulence driving scale, and $v_{\rm A}$ is the Alfvén speed. The fraction $q_{\rm i}$ represents the relative amount of dissipated power transferred to the (non-radiating) ions. [^2] For historic reasons, we refer to $q_{\rm i}$ colloquially as the “heating” fraction, even though $q_{\rm i}$ is defined based on the total power channeled into ions, be it in the form of thermal or nonthermal particles. In a cubic slab with linear size $S$, the compactness can be expressed as $\ell=S^{2}\dot{U}_{\rm esc}\sigma_{\rm T}/(m_{\rm e}c^{3})$ (e.g., Stern et al., 1995). This gives
$\ell\simeq 2\tau_{\rm T}\sigma_{\rm e}(1-q_{\rm i})(v_{\rm A}/c),$ (1)
where $\tau_{\rm T}=\sigma_{\rm T}n_{\rm 0}s_{\rm esc}$ is the Thomson optical depth, $n_{0}$ is the mean electron (or ion) density, $\sigma_{\rm e}=B_{0}^{2}/(4\pi n_{0}m_{\rm e}c^{2})$ is the electron magnetization, and $s_{\rm esc}\simeq S/2$ is the typical distance over which photons need to diffuse in order to escape the source. To obtain the above estimate, we assumed that turbulence is excited on the global scale, such that $s_{0}\simeq s_{\rm esc}$. For simplicity, we will retain this assumption throughout the rest of the paper, but we note that the ratio $s_{\rm esc}/s_{0}$ can be in principle kept as a free parameter of the model. [^3] Our focus on the regime with $s_{\rm esc}\simeq s_{0}$ is also dictated by computational constraints. In particular, at larger values of $s_{\rm esc}/s_{0}$ the simulations become increasingly more challenging because it takes more time to reach a steady state. In principle, the regime with $s_{\rm esc}/s_{0}\gg 1$ could be quite favorable for proton acceleration (see discussion in Sec. II.2). The ion heating fraction $q_{\rm i}$ in (1) is not known a priori; it has been extensively studied in (reduced) kinetic simulations (e.g., Kawazura et al., 2019, 2020; Zhdankin et al., 2019, 2021; Zhdankin, 2021), but so far not in the regime directly applicable to strongly turbulent black-hole coronae. We argue below that $q_{\rm i}$ should be of order unity under conditions expected in black-hole coronae, and we measure $q_{\rm i}\approx 0.6$$0.7$ in simulations (Sec. IV).
Plasma magnetization
Equation (1) essentially constrains the relevant range of plasma magnetizations. Assuming $q_{\rm i}=0.65$, $\tau_{\rm T}=1$, and $m_{\rm i}/m_{\rm e}=1836$ we find that the observationally relevant range $1\lesssim\ell\lesssim 100$ corresponds to:
$0.01\lesssim\sigma_{\rm i}\lesssim 0.2,$ (2)
where $\sigma_{\rm i}=B_{0}^{2}/(4\pi n_{0}m_{\rm i}c^{2})$ is the ion magnetization. Thus, the coronal ions are moderately magnetized whereas the electrons are strongly magnetized, since $\sigma_{\rm e}=\sigma_{\rm i}m_{\rm i}/m_{\rm e}\gg 1$. Interestingly, a recent analysis of global magneto-hydrodynamic (MHD) accretion simulations found $\sigma_{\rm i}\sim 0.1$ in regions identified as the corona (Hankla et al., 2025), which is compatible with our estimate (2). We also note that the collisional coupling between ions and electrons, which is neglected in our model, can become relevant on the low side of the inferred $\sigma_{\rm i}$ range (see Sec. II.4 and Appendix B).
A quantitative example
It may be instructive to put the turbulence parameters into the context of an AGN corona with a given size and black-hole mass. The distance $s_{\rm esc}$ can be taken as a proxy for the typical half-width of the corona (e.g., in spherical geometry this would be the coronal radius $R$). We can estimate
$s_{\rm esc}\sim 10^{13}\!\left(\!\frac{s_{\rm esc}}{10R_{\rm g}}\!\right)\!\left(\!\frac{M}{10^{7}\!\!\;M_{\odot}}\!\right)\,{\rm cm},$ (3)
where $M$ is the black-hole mass, $M_{\odot}$ is the solar mass, and $R_{\rm g}=GM/c^{2}$ is the gravitational radius. The turbulent cascade timescale $t_{0}\sim s_{\rm esc}/v_{\rm A}$ is:
$t_{0}\sim\frac{s_{\rm esc}}{v_{\rm A}}\sim 10^{3}\left(\frac{\sigma_{\rm i}}{0.1}\right)^{\!\!\;\!\!-1/2}\!\!\left(\!\frac{s_{\rm esc}}{10R_{\rm g}}\!\right)\!\!\left(\!\frac{M}{10^{7}\!\!\;M_{\odot}}\!\right)\,{\rm s}.$ (4)
In our model, $s_{\rm esc}/v_{\rm A}$ sets the typical timescale of the short-term coronal variability. Fast X-ray variability, on timescales of a few tens of kiloseconds, has been detected in AGNs (Zhao et al., 2025), and present observations are likely just upper bounds on $t_{0}$. For stellar-mass binary black holes the timescale (4) is of the order of a millisecond, which is consistent with rapid flares detected in Cygnus X-1 (Gierliński & Zdziarski, 2003). Finally, the typical magnetic field strength is
$B_{0}\sim 10^{4}\,\tau_{\rm T}^{1/2}\!\!\;\left(\frac{\sigma_{\rm i}}{0.1}\right)^{\!\!\;\!1/2}\!\left(\!\frac{s_{\rm esc}}{10R_{\rm g}}\!\right)^{\!\!\;\!\!\!\!\;-1/2}\!\!\left(\!\frac{M}{10^{7}\!\!\;M_{\odot}}\!\right)^{\!\!\;\!\!\;\!\!-1/2}\!{\rm G},$ (5)
and the corresponding coronal luminosity $L\simeq(1-q_{\rm i})B_{0}^{2}v_{\rm A}s_{\rm esc}^{2}/6$ (assuming a spherical geometry) is
$L\sim 10^{43}\,\tau_{\rm T}\!\left(\!\frac{1\!-\!q_{\rm i}}{0.35}\!\right)\!\left(\frac{\sigma_{\rm i}}{0.1}\right)^{\!\!\;\!3/2}\!\!\left(\!\frac{s_{\rm esc}}{10R_{\rm g}}\!\right)\!\!\left(\!\frac{M}{10^{7}\!\!\;M_{\odot}}\!\right)\frac{\rm erg}{s}.$ (6)
II.2 Ion Energization
Early theoretical works on the ion heating fraction in kinetic plasma turbulence considered a situation where the turbulent fluctuations introduce only mild perturbations of the particle distribution on top of a thermal background (e.g., Quataert, 1998; Schekochihin et al., 2009; Howes, 2010). The ion plasma beta $\beta_{\rm i}=8\pi n_{0}T_{\rm i}/B_{0}^{2}$ and the ion-electron temperature ratio $T_{\rm i}/T_{\rm e}$ can be then taken as input parameters for calculations of the heating fraction $q_{\rm i}$ (e.g., Kawazura et al., 2019, 2020; Adkins et al., 2025). However, the circumstances are different in open and strongly turbulent domains with $\delta B/B_{0}\sim 1$. There, the distribution of particles entering the domain is significantly altered within a single dynamic timescale, and a steady state average distribution is established via a balance between the rapid particle energization and diffusive escape. Thus, the ion confinement time in a strongly turbulent corona is an important parameter for the determination of the heating partition. In particular, we demonstrate that the confinement time is rather limited even for the moderate-energy particles near the thermal peak of the distribution, when the turbulence is driven on a scale not much smaller than the system size (see Sec. IV.2). This is further corroborated by a recent analysis of particle transport in high-resolution MHD simulations (Kempski et al., 2025), which shows that moderate-energy particles can efficiently diffuse in space by following the strongly turbulent magnetic field lines.
Steady-state kinetic temperature
Let us consider a situation where the ions are nominally in temperature equilibrium with electrons, such that $T_{\rm i}=T_{\rm e}$. Since we focus on large-amplitude turbulence, the typical amplitude of turbulent motions is $\delta v\simeq v_{\rm A}$. For ions with non-relativistic temperatures (to be justified below), the Alfvén speed can be defined based on the cold ion magnetization as $v_{\rm A}=c[\sigma_{\rm i}/(1+\sigma_{\rm i})]^{1/2}$. Note that $v_{\rm A}\simeq c\sigma_{\rm i}^{1/2}$ is a good approximation over the relevant range (see Eq. (2)). The typical turbulent Alfvénic Mach number is then $M_{\rm A}=\delta v/v_{\rm A}\sim 1$, and the sonic Mach number is $M_{\rm s}=\delta v/c_{\rm s}\gg 1$ for typical coronal temperatures of $\sim 100$ keV, where $c_{\rm s}=[(\Gamma_{\rm i}T_{\rm i}+\Gamma_{\rm e}T_{\rm e})/m_{\rm i}]^{1/2}$ is the sound speed and $\Gamma_{s}$ is the species adiabatic index. We therefore expect formation of shocks, which reprocess and heat the plasma. [^4] It may seem that $M_{\rm A}\sim 1$ is a significant limitation for the formation of magnetized shocks. However, when $\delta B/B_{0}\sim 1$ the magnetic field magnitude $|B|$ exhibits order unity fluctuations, and so does the local Alfvén speed. As shown in Sec. IV.1, this enables the formation of regions in space where $M_{\rm A}\sim$ a few, into which the shocks can spread. The electron heating will be offset by their rapid radiative cooling. On the other hand, the ion kinetic temperature will rise until the system reaches a quasi-steady state. Here, we conjecture that the ions are pinned in steady state to the trans-sonic regime with $\delta v\sim v_{\rm A}\sim c_{\rm s}$ and $T_{\rm i}\gg T_{\rm e}$. Since cold plasma is continuously resupplied into the domain, shocks form sporadically in the quasi-steady state. In a trans-sonic and trans-Alfvénic flow, the average ion kinetic temperature and ion-electron temperature ratio are given by:
$\theta_{\rm i}\simeq\frac{\sigma_{\rm i}}{\Gamma_{\rm i}},$ $\frac{T_{\rm i}}{T_{\rm e}}\simeq\frac{\sigma_{\rm e}}{\Gamma_{\rm i}\theta_{\rm e}},$ (7)
where the ion adiabatic index $\Gamma_{\rm i}\approx 5/3$ and $\theta_{s}=T_{s}/(m_{s}c^{2})$ is the dimensionless species kinetic temperature. The average ion kinetic temperature is therefore non-relativistic or very mildly relativistic, which is confirmed by simulations (Sec. IV). Essentially, the trans-sonic state with $M_{\rm s}\sim 1$ is considered here as an attractor of the dynamics. States with $M_{\rm s}\gg 1$ are circumvented by rapid shock heating, while $M_{\rm s}\ll 1$ is prevented by the limited ion confinement time (of the order of $\sim S/v_{\rm A}$ in our model). We emphasize that in the steady state shocks need not be the main energization mechanism, since on average the flow is only trans-sonic. Other dissipation channels can also operate.
Heating fraction
Following the above arguments, $q_{\rm i}$ corresponds to the fraction of turbulence power required to maintain the steady-state ion kinetic temperature given by (7). To obtain $q_{\rm i}$, we assume that ions enter the corona via turbulent advection from a cold external medium, so that their initial energy is roughly the kinetic energy of the turbulent $E\times B$ drift motion; this situation closely resembles our numerical model (Sec. III). The mean ion energy gain is then comparable to their steady state internal energy. Balancing the turbulence power input with the rate of energy gain we obtain: $(3/2)n_{0}T_{\rm i}/t_{\rm conf}\simeq q_{\rm i}B_{0}^{2}v_{\rm A}/(8\pi s_{0})$, where $t_{\rm conf}$ is the average confinement time or lifetime of ions in the corona. In large-amplitude turbulence driven on the box scale, the expected confinement time is of the order of $S/v_{\rm A}$ (Gorbunov et al., 2025). More specifically, in our simulations with $s_{0}\simeq s_{\rm esc}=S/2$ we measure $t_{\rm conf}\simeq 2S/v_{\rm A}$ for a range of $\sigma_{\rm i}$ (see Sec. IV.2). Using this result together with (7) we can estimate $q_{\rm i}$ from the energy balance equation as
$q_{\rm i}\simeq\frac{3}{4\Gamma_{\rm i}}\simeq 0.5.$ (8)
Therefore, we expect that the escaping ions carry away an order-unity fraction of the dissipated power. This is confirmed by our simulations presented in Sec. IV.
Nonthermal acceleration
Above the thermal peak of the distribution (see Eq. (7)), the ion energy spectrum may exhibit a nonthermal tail of the form ${\rm d}N/{\rm d}E\propto E^{-p}$, where $p$ is the power-law index. In the energy range above the thermal peak but below TeV band, a hard nonthermal proton slope (with $1\lesssim p\lesssim 2$) is needed to match the observed neutrino luminosities (Murase et al., 2020; Fiorillo et al., 2024; Mbarek et al., 2024; Lemoine & Rieger, 2025; Saurenhaus et al., 2025), since it is this range that sets the available amount of power for the neutrino-producing TeV protons. Predicting neutrino emission from kinetic plasma simulations requires extrapolations to domains much larger than those accessible to present-day computing resources, which introduces a major uncertainty. In our simulations (Sec. IV), we find that ions develop broad nonthermal spectra, reaching the system-size (Hillas) limit of the simulation box, but the obtained power-law slopes are relatively steep with $p\gtrsim 3$. Macroscopically large domains, particularly those with turbulence driving on scales smaller than the system size, may present a more favorable situation, since in this case the typical ion escape time can become much longer than their acceleration time. However, with larger amounts of energy accumulated in high-energy particles, the properties of the turbulent flow could be modified via the cosmic-ray feedback (Lemoine et al., 2024). Similarly as for the thermal particles, we may therefore expect that some form of self-regulation occurs also for the cosmic-ray component (e.g., as suggested by Lemoine & Rieger, 2025). In addition, nonthermal ion acceleration can be also facilitated by shocks (Inoue et al., 2019). In particular, acceleration by multiple colliding shocks (e.g., Bykov et al., 2013) within a large turbulent volume could provide a possible mechanism for producing a hard ion power law. A more detailed investigation of these aspects requires dedicated simulations with larger boxes, which are left for the future.
II.3 Comptonized Emission Spectrum
X-ray spectrum
Low-energy seed photons introduced into the domain are Comptonized by the coronal electrons and form a power-law continuum with an X-ray photon index $\Gamma_{\!\rm x}$, defined via ${\rm d}N/{\rm d}E\propto E^{-\Gamma_{\!\rm x}}$. The power law extends from the injection energy $E_{0}$ to the peak at $E_{\rm peak}\sim 100$ keV. In steady state, the energy balance between escaping photons and the power received by turbulence reads $(E_{\rm esc}-E_{0})\dot{n}_{\rm ph}\simeq(1-q_{\rm i})B_{0}^{2}v_{\rm A}/(8\pi s_{0})$, where $\dot{n}_{\rm ph}$ is the rate of soft photon supply (equal to the rate of escape in steady state) and $E_{\rm esc}$ is the typical energy of escaping photons. This can be rearranged into
$A-1\simeq\frac{\sigma_{\rm e}(1-q_{\rm i})m_{\rm e}c^{2}}{E_{0}\eta_{\rm ph}},$ (9)
where $A=E_{\rm esc}/E_{0}$ is the photon amplification factor and $\eta_{\rm ph}=\dot{n}_{\rm ph}S/(n_{0}v_{\rm A})$ is an appropriately normalized photon injection rate. Eq. (9) can be alternatively written as $A-1\simeq\ell/\ell_{0}$ (Haardt & Maraschi, 1993), where $\ell_{0}=S^{2}E_{0}\dot{n}_{\rm ph}\sigma_{\rm T}/(m_{\rm e}c^{3})$ is the soft photon compactness supplied to the corona, and the compactness $\ell$ due to turbulent dissipation is estimated in Eq. (1). Beloborodov (1999) obtained a convenient fitting formula of the form $\Gamma_{\!\rm x}\simeq(7/3)(A-1)^{-\delta}$. For AGNs the parameter $\delta\simeq 0.1$. We can then estimate the expected spectral index of the Comptonized radiation as:
$\Gamma_{\!\rm x}\simeq(7/3)\left[\sigma_{\rm e}(1-q_{\rm i})m_{\rm e}c^{2}/(E_{0}\eta_{\rm ph})\right]^{-0.1}.$ (10)
The peak of the spectral energy density can be identified with the mean energy of the coronal electrons. A crude estimate can be obtained by balancing the radiative cooling power with the escaping luminosity. Neglecting Klein-Nishina corrections (Moderski et al., 2005), one can estimate $\Theta_{\rm eff}\sim E_{\rm esc}/[4E_{\rm ph}\tau_{\rm T}(1+\tau_{\rm T})]$, where $E_{\rm ph}\sim E_{\rm esc}$ is the mean energy of photons contained in the source and $\Theta_{\rm eff}=\langle u^{2}/3\rangle$ is an effective dimensionless electron temperature (Grošelj et al., 2024), obtained by averaging the particle 4-velocity $u=\gamma v/c$. The expected peak of the Comptonized spectrum should be then near $E_{\rm peak}\sim m_{\rm e}c^{2}[(1+3\Theta_{\rm eff})^{1/2}-1]$.
MeV tail
Beyond the peak near $\sim$ 100 keV, the spectrum may exhibit an MeV tail, which is shaped by nonthermal electrons (Ghisellini et al., 1993; Zdziarski et al., 1993; Veledina et al., 2011; Zdziarski et al., 2017; Fabian et al., 2017). Nonthermal features are most sensitive to the dissipative plasma physics, and therefore the MeV tail (or lack thereof) can provide important physical constraints. The coronal electrons experience rapid radiative cooling on a timescale shorter than the dynamic time of the turbulent cascade (Grošelj et al., 2024). A very rapid and efficient energization mechanism is required to generate a nonthermal electron tail under such conditions.
In our simulations (Sec. IV) we find that electron acceleration is enabled by collisionless reconnection at intense current sheets, which are embedded intermittently into the turbulent flow (Comisso & Sironi, 2019). Under the expected coronal conditions, reconnection of intense current sheets proceeds in the trans-relativistic and moderate-guide-field regime with $\sigma_{\rm e}\gg 1$, $\sigma_{\rm i}\sim 0.1$, and $\delta B\sim B_{0}$ (Rowan et al., 2017; Werner et al., 2018; Ball et al., 2018; Comisso, 2024). [^5] Existing intermittency models of magnetized plasma turbulence imply that the most intense structures of various sizes roughly preserve the driving-scale fluctuation amplitude (e.g., Chandran et al., 2015). Thus, when $\delta B/B_{0}\sim 1$ at the driving scale, it is reasonable to assume that the most intense current sheets feature a moderate guide field. Note that while the filling fraction of intense sheets may be small, the amount of plasma reprocessed by such sheets during their lifetime can amount to a sizable fraction of the volume (Comisso & Sironi, 2019). For these parameters, Comisso (2024) reports a broken electron power law with a hard exponent $p\approx 1$ between the thermal (i.e., Maxwellian) peak of the distribution and the break at $E_{\rm break}\simeq 0.1\sigma_{\rm e}m_{\rm e}c^{2}$, which is followed by a steeper power law with $p\approx 3.5$ up to some cutoff energy $E_{\rm max}$. An upper bound on $E_{\rm max}$ is provided by radiative cooling. By balancing the radiation drag with acceleration by the reconnecting electric field we find for $\tau_{\rm T}\simeq 1$ that (e.g., see Werner et al., 2019; Sironi & Beloborodov, 2020):
$E_{\rm max}\lesssim E_{\rm rad}\simeq m_{\rm e}c^{2}\left(\frac{\beta_{\rm rec}\sigma_{\rm e}^{1/2}s_{\rm esc}/d_{\rm e}}{f_{\rm KN}\ell}\right)^{1/2},$ (11)
where $d_{\rm e}=[m_{\rm e}c^{2}/(4\pi n_{0}e^{2})]^{1/2}$ is the electron skin depth, $\beta_{\rm rec}\simeq 0.1$ is the collisionless reconnection rate, and $f_{\rm KN}$ is a Klein-Nishina correction factor (Mehlhaff et al., 2021). For a realistic size of an AGN corona (of the order of $\sim 10^{13}d_{\rm e}$ for a $10^{7}$ solar-mass black hole), it can be easily checked that $E_{\rm rad}\gg\sigma_{\rm e}m_{\rm e}c^{2}$, and so the electrons can be accelerated to high energies, perhaps limited only by the size of the (turbulent) reconnection layers (Zhang et al., 2021). The addition of synchrotron loses does not change this conclusion. An expression equivalent to (11) can be obtained for synchrotron loses if one omits $f_{\rm KN}$ and replaces $\ell$ with the magnetic compactness $\ell_{B}\sim\tau_{\rm T}\sigma_{\rm e}$ (Beloborodov, 2017), which is of similar order as $\ell$ (by comparison with Eq. (1)).
In a turbulent box, the expected average nonthermal spectrum is a composite between the one injected at current sheets and the one from electrons in the bulk of the turbulent volume. Outside the current sheets electrons are energized less efficiently, and so their box-averaged spectrum will be generally softer than the one produced at current sheets alone. In particular, it is reasonable to expect that the hard power-law range (below $E_{\rm break}$) may not show up clearly in the average spectrum.
The photons can be in principle upscattered up to the maximum electron energy $E_{\rm max}$. However, in the gamma-ray band the corona becomes opaque to pair production (Murase et al., 2020). The optical depth for absorption of gamma-rays can be estimated as $\tau_{\gamma\gamma}\simeq\tau_{\rm T}n_{1}/(5n_{0})$, where $n_{1}$ is the density of target photons with typical energy $E_{1}\simeq 2m_{\rm e}^{2}c^{4}/E_{2}$, for a given gamma-ray with energy $E_{2}$. For target photons within the range of the X-ray power law, $n_{1}/n_{0}\simeq(n_{\rm ph}/n_{0})(\Gamma_{\rm x}-1)(E_{1}/E_{0})^{1-\Gamma_{\rm x}}$, where $n_{\rm ph}$ is the total photon density. The average ratio $n_{\rm ph}/n_{0}$ can be estimated from (9) by noting that $\dot{n}_{\rm ph}\simeq n_{\rm ph}c/[(\tau_{\rm T}+1)s_{\rm esc}]$. This gives:
$\frac{n_{\rm ph}}{n_{0}}\simeq\frac{\sigma_{\rm e}(v_{\rm A}/c)(1-q_{\rm i})(\tau_{\rm T}+1)m_{\rm e}c^{2}}{2E_{0}(A-1)}.$ (12)
For representative parameters, the ratio $n_{\rm ph}/n_{0}$ is a very large number, of the order of a few ten thousand. This implies that the corona becomes opaque as soon as the gamma-rays can start pair producing with target photons from the peak of the X-ray spectrum, right below the steep MeV tail. A crude but relatively robust estimate for the absorbed gamma-ray photon range is thus given by $E_{\rm abs}\gtrsim 2m_{\rm e}^{2}c^{4}/E_{\rm peak}\sim 5$ MeV.
II.4 Range of Applicability
Our model assumes a strongly turbulent electron-ion corona with weak collisional coupling between ions and electrons. Frequent Coulomb collisions between ions and electrons or copious pair production in MeV-photon collisions can invalidate our assumptions. In Appendix B we estimate the relevant collision rates and find that the applicable $\sigma_{\rm i}$ range is bounded from below by electron-ion collisions and from above by pair production. For representative parameters, we estimate that the turbulent corona can maintain a two-temperature state when $\sigma_{\rm i}\gtrsim 0.04$. According to (1), this corresponds to $\ell\gtrsim 10$. Thus, a two-temperature weakly collisional corona is roughly equivalent to a radiatively compact corona, according to our model. Regarding pair production, we estimate that significant pair enrichment occurs when $\sigma_{\rm i}\gtrsim 0.24$, which corresponds to the high end of the relevant range where $\ell\sim 100$ (see Eq. (2)).
II.5 Alternative Dissipation Scenarios
Alongside turbulence, dissipation in the corona can be mediated by shocks and/or reconnecting current sheets. We stress that our model does not exclude these alternative dissipation mechanisms, and in fact it combines elements of both shocks and reconnecting sheets, which can form naturally within the turbulent flow. A reconnection model was developed by Beloborodov (2017) and explored in a series of numerical simulations (Sironi & Beloborodov, 2020; Sridhar et al., 2021, 2023, 2025). While current sheets can form naturally as part of the turbulent cascade, our present model differs from the work by Beloborodov (2017) in that we consider a more volume-filling dissipation scenario, rather than an isolated reconnecting current sheet in a quiescent background plasma. The latter would imply that dissipation takes place only in a small volume fraction of the corona, which would modify our estimates. Furthermore, the dynamic timescale for the turbulence model is $\sim s_{\rm esc}/v_{\rm A}$, while in the model by Beloborodov (2017) the dynamic timescale is longer by a factor of $\sim 1/\beta_{\rm rec}\sim 10$.
III Numerical Method
In support of our theoretical model, we perform radiative kinetic simulations of driven electron-ion plasma turbulence using the PIC code Tristan-MP v2 (Hakobyan et al., 2023). Our numerical experiments model a local patch of the turbulent corona as an open system for photons and charged particles. Low-energy particles (photons, electrons and ions), representing the influx from a cold medium in thermal equilibrium (e.g., a thin accretion disk), are continuously introduced into the domain. The inserted particles are then energized by the turbulent cascade and leave the box via diffusive escape. The evolution of the radiation in the box and its feedback on the kinetic plasma is determined via a Monte Carlo model of Compton scattering between the electrons and photons (e.g., see Del Gaudio et al., 2020; Grošelj et al., 2024). The heavier ions are not coupled to the radiation; their steady state is thus obtained by a balance between turbulent energization and diffusive escape (Gorbunov et al., 2025). The latter is crucial for overcoming the limitations of closed domains, where ions heat up indefinitely (e.g., Zhdankin et al., 2021).
For computational convenience, we perform the simulations in a 2D periodic box of size $S\times S$. A mean magnetic field $\mathbf{B}_{0}=B_{0}\hat{\mathbf{z}}$ is imposed in the out-of-plane $z$-direction. Turbulence is driven by applying an external time-varying electric current in the form of a “Langevin antenna” (TenBarge et al., 2014). The antenna is configured to drive large-amplitude magnetic fluctuations with $\delta B\sim B_{0}$ and polarization perpendicular to $\mathbf{B}_{0}$. We drive the external current at modes with wavenumbers $(k_{x}S/2\pi,\,k_{y}S/2\pi)=(1,0),\,(0,1),\,(1,1),\,(1,-1)$ and frequencies $\omega_{0}=\pm 1.3v_{\rm A}(2\pi/S)$. The antenna decorrelation rate (TenBarge et al., 2014) is $\gamma_{0}=0.5\omega_{0}$.
The particles are inserted into the domain from an external medium at fixed temperature $T_{0}$. The momenta of the inserted particles are sampled from a Maxwellian distribution for ions and electrons, and from a Planck spectrum for the photons. The thermal momenta of the charged particles are sampled in the local $E\times B$ fluid frame and then boosted into the simulation frame. In order to conserve charge, we insert new ions and electrons at any given time at the exact locations where particles of the same charge just escaped from the box. The mean density $n_{0}$ of ions or electrons is thus fixed by construction. For the seed photons, we prescribe a fixed volumetric injection rate, which is self-consistently balanced by diffusive escape once a steady state has been reached. Following Grošelj et al. (2024) and Gorbunov et al. (2025), we implement escape by tracking the displacement of each particle from its injection location. Particles escape when they diffuse over a distance $s_{\rm esc}=S/2$ in $x$ or $y$, which mimics escape from a turbulent accelerator of size $S$.
We perform a set of 2D simulations at a reduced ion-electron mass ratio of $m_{\rm i}/m_{\rm e}=144$ for different values of the ion magnetization $\sigma_{\rm i}=0.035,\,0.1,\,0.19$. The electron magnetization $\sigma_{\rm e}=\sigma_{\rm i}m_{\rm i}/m_{\rm e}$ is well above unity, as expected for black-hole coronae (see Sec. II.1). The electron skin depth in units of the cell size is $d_{\rm e}=2.5\Delta x$. Compton scattering is resolved on spatial tiles spanning $15$ cells of the PIC grid in each dimension. The time step of the PIC scheme and of the Compton scattering algorithm is $\Delta t=0.29\Delta x/c$. Our square domain has a linear size $S=256\,d_{\rm i}$ in runs with $\sigma_{\rm i}=0.035,\,0.1$ and $S=448\,d_{\rm i}$ in the run with $\sigma_{\rm i}=0.19$, where $d_{\rm i}=(m_{\rm i}/m_{\rm e})^{1/2}d_{\rm e}$ is the ion skin depth. This amounts to a numerical resolution of 7,6802 in runs with $\sigma_{\rm i}=0.035,\,0.1$ and 13,4402 in the $\sigma_{\rm i}=0.19$ run. In all simulations, the average steady-state number of photons per cell (of the PIC grid) is roughly 100. Ions and electrons are each represented with $10$ particles per cell for $\sigma_{\rm i}=0.035,\,0.1$ and with $5$ particles per cell for $\sigma_{\rm i}=0.19$. The external medium in thermal equilibrium, from which we sample the newly inserted particles, has a temperature $T_{0}=10^{-5}m_{\rm e}c^{2}$. The corresponding mean energy of the blackbody seed photons is $E_{0}\approx 2.7\!\!\;\times\!\!\;10^{-5}m_{\rm e}c^{2}$. The reference Thomson optical depth is $\tau_{\rm T}=\sigma_{\rm T}n_{0}s_{\rm esc}=1$. For each run, we choose a suitable soft-photon injection rate $\eta_{\rm ph}=\dot{n}_{\rm ph}S/(n_{0}v_{\rm A})$, so that the obtained X-ray spectral index is in good agreement with observations (see Eq. (10)). To this end, we performed a series of preparatory test runs, until an appropriate choice of $\eta_{\rm ph}$ was found for a given $\sigma_{\rm i}$. In our production runs, we use $\eta_{\rm ph}=3.18\;\!\!\times\;\!\!10^{3},\,9.07\!\!\;\times\!\!\;10^{3},\,2.10\!\!\;\times\!\!\;10^{4}$ for $\sigma_{\rm i}=0.035,\,0.1,\,0.19$, respectively.
The predicted spectra are compared with X-ray observations of NGC 4151 from Swift/BAT, INTEGRAL, and NuSTAR. The Swift/BAT and INTEGRAL data represent long-term averages over several years of observations. The NuSTAR data are from Nov 14, 2012. The PIC simulation spectra are corrected for absorption and reflection when comparing with observations. We adopt a simple model, consistent with an earlier study of the same source (Hinkle & Mushotzky, 2021). For the reflection component, use the reflect model (Magdziarz & Zdziarski, 1995) of xspec (Arnaud, 1996) with a reflection factor $R=0.494$ and inclination angle $i=45^{\circ}$. The spectra are additionally corrected for absorption using the ztbabs model of xspec, with hydrogen column density $N_{\rm H}=10.3\!\!\;\times\!\!\;10^{22}/{\rm cm}^{2}$, redshift $z=0.003$, and elemental abundances from Wilms et al. (2000).
IV Simulation Results
Below we present results from our set of local PIC simulations of strongly turbulent electron-ion coronae of accreting black holes. The simulations are performed for different strengths of the ion magnetization $\sigma_{\rm i}$, which correspond to different values of the steady-state radiative compactness $\ell$. In Sec. IV.1 we investigate the properties of the turbulent flow, which is followed by the analysis of charged particle and photon spectra in Sec. IV.2.
IV.1 Properties of the Turbulent Flow
Global evolution
Let us first demonstrate how our simulations settle into a trans-sonic and trans-Alfvénic steady state, consistent our theoretical arguments (Sec. II.2). Fig. 1 shows the time evolution of various box-averaged quantities. A quasi-steady state is attained in roughly 4 Alfvén crossing times $S/v_{\rm A}$. In the bottom two panels we show the evolution of the sonic Mach number $M_{\rm s}=v_{\rm rms}/c_{\rm s}$ and of the Alfvénic Mach number $M_{\rm A}=v_{\rm rms}/v_{\rm A}$, where $v_{\rm rms}$ is the root-mean-square ion bulk velocity and $c_{\rm s}=(\Gamma_{\rm i}T_{\rm i}/m_{\rm i})^{1/2}$ is the ion sound speed (using $\Gamma_{\rm i}=5/3$ for the adiabatic index). We find that the steady state is characterized by $M_{\rm A}\sim M_{\rm s}\sim 1$. For the considered range of ion magnetizations ($\sigma_{\rm i}\lesssim 0.2$), the corresponding steady-state proper ion kinetic temperature $T_{\rm i}$ is non-relativistic or very mildly relativistic, as anticipated in Eq. (7). The box-averaged proper electron kinetic temperature is of the order of $T_{\rm e}\sim 0.1m_{\rm e}c^{2}$, and shows a mild decline with increasing $\sigma_{\rm i}$. The latter can be attributed to the increased cooling strength at larger values of the compactness (Grošelj et al., 2024), which scales as $\ell\sim\sigma_{\rm e}(v_{\rm A}/c)\sim(m_{\rm i}/m_{\rm e})\sigma_{\rm i}^{3/2}$ (see Eq. (1)). In the top panel of Fig. 1 we calculate $\ell=S^{2}\dot{U}_{\rm esc}\tau_{\rm T}/(s_{\rm esc}n_{0}m_{\rm e}c^{3})$ by measuring the box-averaged power of escaping photons per unit volume. In steady state, we find $\ell\approx 0.6,\,2.5,\,8.5$ in simulations with $\sigma_{\rm i}=0.035,\,0.1,\,0.19$, respectively. The run with $\sigma_{\rm i}=0.19$ is therefore most representative of the high-compactness regime. Note that the obtained $\ell$ values are lower than those expected for a proton-electron plasma due to our use of a reduced ion-electron mass ratio ($m_{\rm i}/m_{\rm e}=144$), which is a choice imposed by computational constraints.
Figure 1: Time evolution and approach to steady state in our radiative PIC simulations of turbulence with ion magnetizations $\sigma_{\rm i}=0.035,\,0.1,\,0.19$. Shown from top to bottom are the box-averaged compactness $\ell$, proper electron kinetic temperature $T_{\rm e}$, proper ion kinetic temperature $T_{\rm i}$, the turbulent sonic Mach number $M_{\rm s}$, and the Alfvénic Mach number $M_{\rm A}$.
Figure 2: Visualization of turbulent fields in our simulation with $\sigma_{\rm i}=0.19$ at time $t=6.86\,S/v_{\rm A}$. The left panels show the proper ion and electron kinetic temperatures ($T_{\rm i}$ and $T_{\rm e}$). In the middle panels we show the magnetic field magnitude $B$ and the out-of-plane electric current $J_{z}$. Finally, the right panels depict the ion density $n$ and the density of fast-mode fluctuations $n_{\rm fast}$. For easier comparison with the total density $n$, we include the mean value in the visualization of the fast mode (i.e., $n_{\rm fast}=\delta n_{\rm fast}+n_{0}$). An animated version of this figure is available online at https://youtu.be/0W-b242WMhw. The animation lasts 50 s and shows the spatiotemporal evolution of the turbulent fields from the start ($tv_{\rm A}/S=0$) to the end ($tv_{\rm A}/S=10.3$) of the simulation, in the same format as the static figure.
Spatial structure
In Fig. 2 we depict the spatial structure of turbulent fluctuations in the quasi-steady state of the simulation with $\sigma_{\rm i}=0.19$. It can be seen that the plasma develops a two-temperature state with $T_{\rm i}\gg T_{\rm e}$. The turbulent field is populated with intense current sheets and magnetized shocks, which energize the particles. By means of an MHD mode decomposition (see Appendix C), we extract the density fluctuations corresponding to fast modes and find that shocks form via steepening of large-amplitude fast waves. The presence of fast modes is a natural consequence of our turbulence driving with $\delta B/B_{0}\sim 1$. [^6] A second-order expansion of the magnetic field magnitude gives $B/B_{0}\approx 1+\delta B_{z}/B_{0}+(\delta B_{\perp}/B_{0})^{2}/2$, where $\delta B_{z}$ and $\delta B_{\perp}$ are the fluctuations parallel and perpendicular to $\mathbf{B}_{0}$, respectively. Thus, even though our forcing induces Alfvénic-like fluctuations with $\delta B\approx\delta B_{\perp}$, at large amplitudes a portion of the injected power is converted into compressible fast modes via magnetic-pressure fluctuations. Unlike the rest of the turbulent fluctuations (e.g., the plasma density $n$ in Fig. 2 modulo the fast mode contribution $n_{\rm fast}$, or the $J_{z}$ fluctuations), which develop a filamentary structure typical for Alfvénic turbulence, the fast modes feature a combination of relatively smooth large-scale structures and shock discontinuities (seen most clearly in snapshots of $B$ and $n_{\rm fast}$). Snapshots of $B$ also show that shocks tend to spread into regions with below-average magnetic field magnitude ($B<B_{0}$), where the local Alfvén speed is lower than the reference value $v_{\rm A}\simeq c\sigma_{\rm i}^{1/2}$. This allows for formation of shocks with Alfvénic Mach numbers up to $\sim$ a few. For a more detailed analysis of shock properties see Appendix A.
Figure 2 also shows that the fluctuations in the ion and electron temperatures are correlated with the turbulent structures. In particular, the electron kinetic temperature is well correlated with electric current sheets. On the other hand, the ion temperature is enhanced at both shocks and current sheets, but the apparent correlation is weaker than for electrons. This is because the ions do not experience the rapid radiative cooling, and so their temperature fluctuations can spread over large regions of the volume away from the local energization sites. In Appendix A we analyze the local particle distributions near current sheets and shocks and find that electrons are injected into nonthermal distributions predominantly at current sheets, whereas the ions can be injected at both shocks and current sheets. We also find that the nonthermal electrons develop anisotropic distributions with small pitch angles with respect to the local magnetic field (on this note, see also Comisso, 2024; Vega et al., 2024).
Spectrum of turbulence
In Fig. 3 we analyze the 1D turbulent energy spectra ${\mathcal{E}}(k_{\perp})$ as a function of the wavenumber $k_{\perp}$ perpendicular to $\mathbf{B}_{0}$. Over the MHD scale range ($k_{\perp}\lesssim 1/d_{\rm i}$) the turbulence spectrum shown in the bottom panel of Fig. 3 forms a power law with a slope close to 3/2-3/2. This is consistent with refined models of MHD turbulence that incorporate the dynamic alignment effect (Boldyrev, 2006). Near the transition into the kinetic range the spectrum then steepens as a result of dispersive wave physics and collisionless damping of the turbulent cascade (Howes, 2015). The position of the spectral break is more closely aligned with the scale of the ion thermal gyroradius $\rho_{\rm i}=\sqrt{\beta_{\rm i}}\,d_{\rm i}$ (estimated from the steady-state ion beta $\beta_{\rm i}\approx 0.57$), rather than with $d_{\rm i}$, in accordance with theoretical expectations for low-frequency Alfvénic turbulence (Schekochihin et al., 2009). The position of the break could be also affected by shocks, which have a typical width of a few ion skin depths (Appendix A). From the steady-state spectrum we also estimate the turbulence correlation length, which can be taken as a proxy for the driving scale $s_{0}$. We find $s_{0}\simeq\pi\left(\int k_{\perp}^{\;\!-1}{\mathcal{E}}(k_{\perp}){\rm d}k_{\perp}\right)/\left(\int{\mathcal{E}}(k_{\perp}){\rm d}k_{\perp}\right)\approx S/3$, and so the effective ratio $s_{\rm esc}/s_{0}\approx 3/2$ is slightly larger than unity for our choice of the forcing.
Figure 3: Turbulent energy spectra in our simulation with $\sigma_{\rm i}=0.19$, averaged over the steady state from $t=4\,S/v_{\rm A}$ until the end of the run at $t=10.3\,S/v_{\rm A}$. In the bottom panel we show the 1D energy spectrum ${\mathcal{E}}(k_{\perp})$ as a function of the perpendicular wavenumber $k_{\perp}$. Power-law slopes are shown for reference. In the top panel we show the relative energy content of fluctuations identified as the MHD Alfvén, slow, and fast mode (see Appendix C for details).
3D radiative PIC simulations of pair plasmas by Grošelj et al. (2024) demonstrated that turbulent bulk motions can be subject to radiative damping, which gives rise to steeper (non-universal) turbulence spectra over the MHD scale range. On the other hand, the spectrum shown in Fig. 3 does not show any apparent signs of abnormal steepening, and is instead reminiscent of turbulent cascades in more conventional weakly radiating plasmas, such as the solar wind (Kiyani et al., 2015). To put our results in the context of the previous work by Grošelj et al. (2024), we estimate the amount of power passed to the photons via bulk Comptonization. Presently we are unaware of a generally accepted definition of bulk Comptonization in a kinetic electron-ion plasma. Here, we employ a definition based on the transfer of bulk momentum between electrons and photons, which seems like natural choice from an energetics point of view. At each time step, we compute the local net change of the photon momentum $\delta\mathbf{P}_{\!\rm tile}$, and from there we estimate $\Delta E_{\rm bulk}=\sum c\left(|\mathbf{P}_{\!\rm tile}+\delta\mathbf{P}_{\!\rm tile}|-|\mathbf{P}_{\!\rm tile}|\right)$, where the sum runs over the spatial tiles of our Monte Carlo Compton scattering scheme (see Sec. III), and $\mathbf{P}_{\rm tile}$ is the net photon momentum in the tile before scattering. Each tile contains of the order of $\sim 10^{4}$ computational photons and $\sim 10^{3}$ electrons, so that $\delta\mathbf{P}_{\!\rm tile}$ is a measure of the local bulk momentum transfer between the electron and photon gas. The power lost via bulk Comptonization is then obtained by averaging $\Delta E_{\rm bulk}/\Delta t$ over steady state. This gives for the bulk Comptonization fraction $f_{\rm bulk}\approx 0.18,\,0.37,\,0.51$ in simulations with $\sigma_{\rm i}=0.035,\,0.1,\,0.19$, respectively. That $f_{\rm bulk}$ increases with $\sigma_{\rm i}$ is understandable, since the magnetization sets the Alfvén speed, and in turn the typical magnitude of the turbulent bulk motions. This suggests that the turbulent motions in an electron-ion corona with $\sigma_{\rm i}\sim 0.1$ may not be sufficiently relativistic to give rise to a significant radiative steepening of the turbulence spectrum. In this regard, it may seem surprising that the measured values of $f_{\rm bulk}$ are still relatively high. The bulk momentum of a (mildly) relativistic fluid is $w\Gamma_{\rm bulk}^{2}\mathbf{v}_{\rm bulk}/c^{2}$, where $w$ is the proper enthalpy density, $\mathbf{v}_{\rm bulk}$ is the bulk velocity and $\Gamma_{\rm bulk}=\left[1-(v_{\rm bulk}/c)^{2}\right]^{-1/2}$. Thus, the amount of power lost via bulk Comptonization can be enhanced by mildly relativistic electron temperatures (via the enthalpy term), according to our procedure for estimating $f_{\rm bulk}$. This may be a reason why the estimated bulk Comptonization fractions are relatively high.
To further analyze the nature of turbulent fluctuations, we decompose in the top panel of Fig. 3 the spectra into contributions from the MHD Alfvén, slow, and fast modes (see Appendix C for details). At the largest scales ($k_{\perp}\lesssim 0.05/d_{\rm i}$), fast modes contain roughly 30% of the total energy, less than 10% is contained in slow modes, and the rest is carried by Alfvén modes. The sum of fluctuations identified as either the Alfvén or slow mode is relatively stable over the MHD scale range ($k_{\perp}\lesssim 1/d_{\rm i}$), in contrast to the fluctuations identified as Alfvén waves only, which feature an apparent decline with growing $k_{\perp}$. According to our classification scheme (see relation (C2) in Appendix C), any pressure balanced fluctuation is categorized as the slow mode. However, in a kinetic plasma Alfvén waves gradually develop a finite compressibility as their wavelength approaches the ion kinetic scales (Howes et al., 2006). In addition, slow modes are not expected to dynamically couple to the Alfvénic cascade and are instead merely passively mixed by the Alfvénic turbulence (Schekochihin et al., 2009). This suggests that the declining Alfvén mode fraction with increasing $k_{\perp}$ can be largely attributed to the conversion of the Alfvén modes into compressible kinetic Alfvén waves. Therefore, the sum $\mathcal{E}_{\rm Alfven+slow}$ may be a better representation of the true energy content of the Alfvénic cascade, while the true energy content of slow modes is likely minor.
It is known that the cascade of turbulent fast modes is isotropic in full 3D geometry (Cho & Lazarian, 2003; Takamoto & Lazarian, 2016), in contrast to the Alfvénic cascade which is anisotropic in wavenumber space (Goldreich & Sridhar, 1995). Therefore, 3D kinetic simulations are required for a complete picture of how the partitioning among different modes shapes the nature of the turbulent cascade and the resulting particle energization. An exact 3D equivalent of our 2D PIC simulations is computationally infeasible at present. 3D aspects of the problem could be instead studied with reduced-kinetic or fluid-like models, provided that such approximations are appropriate for the specific question at hand.
IV.2 Energy Partitioning, Nonthermal Acceleration and Emission Spectrum
Ion and electron spectra
We now analyze the distribution of particles energized by the turbulent cascade. The steady-state distributions are shown in Fig. 4. The ion spectra are shown in the middle panel and the electron spectra in the bottom panel. Both the ion and the electron distributions feature a quasi-thermal (i.e., Maxwellian) part and a high-energy nonthermal tail. In accordance with Eq. (7), the thermal peak of the ion distribution shifts to higher energies at larger values of $\sigma_{\rm i}$, which correspond to larger values of the radiative compactness $\ell$. On the other hand, the position of the electron thermal peak is independent of $\sigma_{\rm i}$ (or $\ell$), which shows that the effective electron temperature is primarily set by the optical depth (equal to $\tau_{\rm T}=1$ in all runs), as explained in Sec. II.3.
Figure 4 also shows that higher magnetizations lead to more pronounced nonthermal features. With the exception of the low-compactness simulation ($\ell\approx 0.6$), the nonthermal ions are accelerated up to the maximum energy set by the size of the simulation box $E_{\rm max}=m_{\rm i}c^{2}\sigma_{\rm i}^{1/2}s_{\rm esc}/d_{\rm i}$ (dotted vertical lines in the middle panel of Fig. 4), where the particle gyroradius $\rho_{\max}=s_{\rm esc}$. The local power-law index in the nonthermal parts of the ion distribution is $p\gtrsim 3$. The electron power-law index hardens with growing magnetization (or compactness). For $\ell>1$, we find $p\approx 3.5$ for electrons. This is broadly consistent with our theoretical discussion from Sec. II.3, where we argue that electrons are predominantly accelerated at intense reconnecting current sheets, in which case a power-law index not far from $p\approx 3.5$ is expected for $\sigma_{\rm i}\sim 0.1$ and $\delta B/B_{0}\sim 1$ (Comisso, 2024). It can be also noticed that in between the electron thermal peak and the power-law tail lies a suprathermal range, where the spectrum is relatively flat. This could be in part related to the hard power-law range expected for high-$\sigma_{\rm e}$ reconnection (see Sec. II.3), but in a turbulent volume the hard power law could be obscured by contributions from particles outside the current sheets, where the spectrum is softer. For a more detailed analysis of particle injection into nonthermal populations near shocks and current sheets see Appendix A.
Figure 4: Steady-state particle spectra in our simulations with different strengths of the ion magnetization $\sigma_{\rm i}$. The middle panel shows the ion energy spectra and the bottom panel the electron spectra. Energy is measured in units of the particles’ own rest mass ($m_{\rm i}c^{2}$ for ions and $m_{\rm e}c^{2}$ for electrons). Power-law slopes are indicated with black dotted lines for reference. In the bottom panel we fit a Maxwellian distribution to the low-energy part of the spectrum (dashed black curve). The top panel shows the ion escape time $t_{\rm esc,i}$. Finally, the inset shows the ion heating fraction.
In the top panel of Fig. 4 we calculate the energy-dependent ion escape time $t_{\rm esc,i}=f_{\rm i}(E)/\dot{f}_{\rm esc,i}(E)$, where $\dot{f}_{\rm esc,i}(E)$ is the measured flux of escaping ions per unit energy and $f_{\rm i}(E)={\rm d}N_{\rm i}/{\rm d}E$ is the distribution of ions contained in the box (see also Gorbunov et al., 2025). At energies above the thermal peak, the measured escape time is of the order of an Alfvén crossing time $S/v_{\rm A}$. This shows that even moderate-energy particles are not well confined by the source, when the turbulence is strongly driven on a scale comparable to the system size. An alternative energy-independent measure may be obtained by calculating the average lifetime or confinement time of particles in the box, from the moment they are introduced into the domain and until they escape. In steady state, we measure an average confinement time of roughly $2S/v_{\rm A}$, which is in reasonable agreement with the energy-dependent measure shown in the top panel of Fig. 4.
Ion heating fraction
The inset plot of Fig. 4 shows the relative amount of turbulence power deposited into the escaping ions. We calculate the ion fraction as $q_{\rm i}=(L_{\rm i}-L_{\rm i0})/\sum_{s}(L_{s}-L_{s0})$, where $L_{s}$ is the average escaping luminosity of a given species ($s=$ ions, electrons, or photons), and $L_{s0}$ is the average power from new particles supplied into the box (essentially, $L_{s}-L_{s0}$ is the average power species $s$ receives from the turbulent cascade). We measure $q_{\rm i}\approx 0.72,\,0.73,\,0.63$ in simulations with $\sigma_{\rm i}=0.035,\,0.1,\,0.19$, respectively. This is broadly consistent with our theoretical estimate, which predicts $q_{\rm i}\simeq 0.5$ (Eq. (8)). Our theoretical argument is rather generic and does not specify how exactly the ions are energized in the steady state. More detailed heating predictions could be built upon specific energization mechanisms. For representative parameters of the steady state ($\sigma_{\rm i}\sim 0.1,\,\beta_{\rm i}\sim 1,\,\delta B/B_{0}\sim 1,\ T_{\rm i}\gg T_{\rm e}$), heating prescriptions based on either turbulent wave-particle interactions (Howes, 2010; Kawazura et al., 2019, 2020; Adkins et al., 2025) or magnetic reconnection (Rowan et al., 2017; Werner et al., 2018; Rowan et al., 2019) would be compatible with values close to $q_{\rm i}\sim 0.5$, while heating by low Mach number shocks would imply $(1-q_{\rm i})\ll 1$ (Sironi & Tran, 2024). We emphasize that $q_{\rm i}$ measures the total fraction of energy deposited into the escaping ions, in the form of thermal or nonthermal particles. In other words, $q_{\rm i}$ can be taken as an upper limit for the available fraction of energy that can be deposited into cosmic-ray ions. Finally, it is worth mentioning that our lowest magnetization run with $\sigma_{\rm i}=0.035$ would be realistically in the regime with significant collisional coupling between ions and electrons (see Appendix B). The $\sigma_{\rm i}=0.035$ run is included here to better illustrate the general dependence on $\sigma_{\rm i}$ in the absence of ion-electron collisions, but caution is advised when applying the low-$\sigma_{\rm i}$ results to real sources, where the collisional coupling could play a role.
Emission spectrum
Figure 5 shows the predicted emission spectra from our model and compares the results with observations of NGC 4151 from NuSTAR, Swift/BAT, and INTEGRAL. We take the NuSTAR measurements as a reference and normalize the other data by matching to the NuSTAR luminosity in the 26.2 to 68.2 keV band (where NuSTAR spectra overlap with the INTEGRAL and Swift results). The intrinsic steady-state spectra of escaping photons are shown with dashed curves. The solid curves show the PIC simulation spectra corrected for reflection and absorption (see Sec. III for details).
Figure 5: Comparison of predicted emission spectra with X-ray observations of NGC 4151. Dashed curves show the intrinsic steady-state spectra obtained directly from our PIC simulations with different strengths of the radiative compactness $\ell\approx 0.6,\,2.5,\,8.5$. Solid curves show the spectra corrected for absorption and reflection. Shaded bands indicate the ranges over which the simulated spectra vary during the averaging interval from $t\approx 4S/v_{\rm A}$ to $t\approx 10S/v_{\rm A}$.
Our simulated spectra are in excellent agreement with observations of NGC 4151. This demonstrates that our kinetic simulations probe the observationally relevant regime of turbulence, and so the obtained physical insight (e.g., concerning the energy partitioning) can be put into the context of real sources. The intrinsic X-ray photon index is close to $\Gamma_{\rm x}=1.77$ (dotted black line in Fig. 5), which is in reasonable agreement with previous studies of NGC 4151 (e.g., by Lubiński et al., 2016, who found $\Gamma_{\rm x}=1.76$). For our choice of model parameters (see Sec. III) and for the measured values of $q_{\rm i}$, Eq. (10) predicts $\Gamma_{\rm x}\approx 1.76,\,1.77,\,1.75$ in runs with $\sigma_{\rm i}=0.035,\,0.1,\,0.19$, respectively. The peak of the (intrinsic) spectrum is near 80 keV, which is as well in agreement with our theoretical estimate $E_{\rm peak}\sim m_{\rm e}c^{2}\left\{[1+3/(4\tau_{\rm T}(1\!\!\;+\!\!\;\tau_{\rm T}))]^{1/2}-1\right\}\sim 90$ keV. Thus, our numerical results are well aligned with theoretical expectations for the model.
Figure 5 also shows that the predicted X-ray spectrum is degenerate with respect to model parameters, in the sense that simulations employing different values of $\ell$ can be matched to observations. However, the differences become apparent above the peak of the spectrum, where larger values of the radiative compactness lead to a more pronounced MeV tail. As discussed in Sec. II.3, the MeV tail is shaped by nonthermal electrons, which are sensitive to the microphysical details of the dissipation. In particular, larger values of the compactness imply higher electron magnetizations, which give rise to harder nonthermal tails from electrons accelerated at reconnecting current sheets. Moreover, in the range around 1 MeV the spectrum is nearly unaffected by the (parameter-dependent) reflection modeling. Thus, future measurements with MeV-band instruments (Tomsick et al., 2024) can provide important physical constraints on the emission mechanism. Past attempts to measure the MeV tail in NGC 4151 (Zdziarski et al., 1996) suggest that the spectrum indeed extends into the MeV range, in broad agreement with our predictions for $\ell>1$.
V Discussion and Conclusions
We developed a local emission model of black-hole coronae based on dissipation in a strongly turbulent electron-ion plasma. Our theoretical estimates (Sec. II) are informed by a set of radiative PIC simulations (Sec. IV), which include self-consistent Compton scattering, as well as injection and diffusive escape of charged particles and photons. The simulations allow us to track from first principles the turbulent partitioning of energy among ions, electrons, and photons. We show that the turbulent cascade regulates itself into a trans-sonic state, which features sporadic formation of magnetized shocks (Sec. IV.1). In steady state, the ion kinetic temperature far exceeds the electron kinetic temperature ($T_{\rm i}\gg T_{\rm e}$). Complementary to alternative arguments, based e.g. on assuming that the ions are near the virial temperature (e.g., Bambic et al., 2024), we interpret the two-temperature state as a natural outcome of the interplay between the dissipative microphysics and radiation. In order to maintain the expected steady-state ion kinetic temperature, an order-unity fraction of the available power needs to be deposited into the ions (see Sec. II.2). In our simulations, we measure $0.6\lesssim q_{\rm i}\lesssim 0.7$ for the ion heating fraction, which shows that most of the dissipated power is deposited into the ions.
The predicted emission spectra obtained from our PIC simulations are confronted with X-ray observations of NGC 4151, demonstrating excellent agreement (Sec. IV.2). We show that the MeV tail, which is presently poorly constrained by observations, can provide important insight into the nonthermal physics of the coronal electrons. In particular, we find that higher values of the radiative compactness $\ell$ give rise to a more pronounced MeV tail, which is shaped by nonthermal electrons. Our predictions can be tested with future MeV-band instruments such as COSI (Tomsick et al., 2024). Presently, the diffuse cosmic X-ray background gives an upper limit on how strong on average the MeV tail generated by the nonthermal electrons might be. Measurements of the X-ray background indeed show an excess over conventional models in the $\sim$ 150 – 1000 keV band (e.g., see Figure 9 from Marcotulli et al., 2022). In addition, MeV photons can be also generated in electromagnetic cascades of high-energy gamma-rays, born in hadronic reaction chains (e.g., Murase et al., 2020). Modeling of electromagnetic cascades is beyond the scope of the present work.
While our model gives an apparent good match to the observed X-ray spectrum of a bright AGN, it falls short of being able to predict the emission in lower-energy bands, such as the UV/optical (e.g., Steffen et al., 2006; Duras et al., 2020) or the mm band (e.g., Kawamuro et al., 2022; Ricci et al., 2023), and their connection to the X-ray luminosity. We stress that our local model only tracks the soft photons intercepted by the hot corona. Realistically, only a fraction of the UV/optical photons from the disk is actually intercepted by the corona (Done et al., 2007). Similarly, the recently observed mm emission may come from a large-scale outflow launched from the corona (Hankla et al., 2025). Thus, accurate multi-wavelength predictions require a global model for the extended accretion disk and its corona.
Finally, we find that a strongly turbulent corona can efficiently accelerate nonthermal particles (Sec. IV.2). We investigate the origins of nonthermal acceleration and show that the electrons are preferentially injected into nonthermal power laws at intense current sheets, while the ions can be injected at both shocks and current sheets (Appendix A). In our limited-size simulation boxes, the nonthermal ions can be accelerated up to the system size (Hillas) limit, albeit with a relatively steep spectrum (with power-law index $p\gtrsim 3$). Simulations with significantly larger boxes are required to confidently predict the shape of the nonthermal ion (i.e., proton) spectrum up to the neutrino-producing TeV range. However, regardless of the exact shape of the ion spectrum, the maximum fraction of the coronal power deposited into cosmic rays is limited from above by the ion heating fraction $q_{\rm i}$. The relatively large values of the measured heating fraction ($0.6\lesssim q_{\rm i}\lesssim 0.7$) present a quite favorable situation for producing energetically significant amounts of cosmic-ray particles.
In conclusion, we showed that strongly driven and radiative electron-ion turbulence can generate nonthermal ion and electron distributions, while producing at the same time emission spectra consistent with X-ray observations of black-hole coronae. The steady state is a two-temperature trans-sonic flow with ions much hotter than electrons. Present computational constraints prevent us from further improving the realism of our numerical model (e.g., the box size). A promising future direction may be to disentangle the subject into a set of complementary questions, such as the physics of the MeV tail or nonthermal proton acceleration, which can be investigated with simplified models specifically designed to address a particular challenge.
We gratefully acknowledge helpful discussions with F. Bacchini, L. Sironi, R. Mbarek, B. Ripperda, H. Krawczynski, A. Hankla, E. Gorbunov, D. Caprioli, L. Comisso, and A. Levinson. D.G. is supported by the Research Foundation–Flanders (FWO) Senior Postdoctoral Fellowship 12B1424N. This work was supported by NASA (grant 80NSSC22K1054) and the Simons Foundation (grant 00001470; A.P.), and was facilitated by the Multimessenger Plasma Physics Center (MPPC; A.P. and A.B.) under NSF grant No. PHY-2206607. A.P. additionally acknowledges support from an Alfred P. Sloan Fellowship and a Packard Foundation Fellowship in Science and Engineering. The resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the FWO and the Flemish Government. We acknowledge LUMI–BE for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium, through a LUMI–BE Regular Access call. Simulations were additionally performed on NASA Pleiades (GID s2056).
Appendix A Analysis of Shocks and Current Sheets
Figure 6: Structure of a shock inside the turbulent box of the run with $\sigma_{\rm i}=0.19$. Top panels show the fast-mode density and the $y$-component of the ion bulk velocity in a fraction of the domain. 1D shock profiles (middle and bottom panel) are extracted along the path indicated with arrows. In the middle panel we show the total density $n$, the fast-mode density $n_{\rm fast}$, the magnetic field magnitude $B$, and the bulk velocity component parallel to the shock normal. In the bottom we show profiles of the ion and electron kinetic temperature (solid curves), along with expected values for adiabatic compression (dashed curves).
Shock structure
In Fig. 6 we analyze the structure of a representative shock extracted from the turbulent volume of our simulation. In the top panels we visualize the structure of the fast-mode density and the $y$-component of the ion bulk velocity. Along the path indicated by the arrows we extract 1D shock profiles, which are shown in the bottom two panels of Fig. 6. The profiles shown in the middle panel reveal a jump (near $s/d_{\rm i}\approx 30$) in the magnetic field and fluid velocity, which can be associated with the location of the shock. The shock transition is a few ion skin depths wide. The profile of the total density shows the corresponding jump less clearly than the density of fast-mode fluctuations alone (see Appendix C), since the total density includes pressure-balanced fluctuations not belonging to fast modes. [^7] Note that for the pressure-balanced fluctuations $\delta n/n_{0}\simeq-\alpha^{-1}\,\delta B/B_{0}$ (see Appendix C), with $\alpha^{-1}\approx 3.1$ at the time analyzed in Fig. 6. Therefore, the pressure-balanced fluctuations show up more clearly in the profiles of $n$ rather than $B$. Based on the fast-mode profile, we determine a density compression factor $r\approx 3.2$ between the shock upstream ($s/d_{\rm i}<30$) and downstream ($s/d_{\rm i}>30$). The velocity jump across the shock transition is $\delta v\approx 0.53c$. The corresponding Alfvénic Mach number is $M_{\rm A}=\delta v/v_{\rm A,u}\approx 2.3$, where $v_{\rm A,u}$ is the upstream Alfvén speed. The same value is found for the sonic Mach number $M_{\rm s}=\delta v/c_{\rm s,u}$, where $c_{\rm s,u}$ is the upstream sound speed. The shock analyzed in Fig. 6 is quasi-perpendicular; the angle of the upstream $\mathbf{B}$ field relative to the shock normal is $\approx 70^{\circ}$. In general, the shocks embedded into the turbulent box feature a broad range of upstream magnetic field orientations, but the quasi-perpendicular case seems to be the most typical.
Figure 7: Injection of particles into nonthermal populations near shocks and current sheets. On the top left we show the out-of-plane electric current $J_{z}$ and the divergence of the bulk velocity $\nabla\cdot{\boldsymbol{\beta}}_{\rm bulk}$ in a fraction of the 2D simulation domain. On the bottom left we show the density of nonthermal electrons $n_{\rm e,nt}$ and nonthermal ions $n_{\rm i,nt}$. The red and green squares mark the representative locations of a current sheet and a shock, respectively, from which we extract the local particle distributions shown on the bottom right side of the figure. On the top right side we show the mean ion and electron pitch angle relative to $\mathbf{B}$, computed in the local $E\times B$ frame.
In the bottom panel of Fig. 6 we show the profiles of the ion and electron proper kinetic temperature. For reference, we also calculate the profiles corresponding to adiabatic compression $T_{s,{\rm ad}}=T_{s,{\rm u}}(n_{\rm fast}/n_{\rm fast,u})^{\Gamma_{s}-1}$, where the subscript $\rm u$” denotes the mean upstream values and $\Gamma_{s}$ is the species adiabatic index (we use $\Gamma_{s}=5/3$ for both species). The electrons do not show any appreciable temperature increase beyond expectations for adiabatic compression. In fact, the values of $T_{\rm e}$ are even somewhat below $T_{\rm e,ad}$, which may be due to the rapid radiative cooling experienced by the electrons. On the other hand, the downstream ion kinetic temperature raises above the expectations for adiabatic compression, which indicates entropy production across the shock transition.
Particle energization at shocks and current sheets
In Fig. 7 we analyze the local particle distributions at shocks and current sheets in the simulation with $\sigma_{\rm i}=0.19$. Current sheets can be identified as intense sheetlike structures in the out-of-plane current density $J_{z}$, while shocks can be seen as near-discontinuities in the divergence of the ion bulk velocity $\nabla\cdot{\boldsymbol{\beta}}_{\rm bulk}$ (top left side of Fig. 7). Using the spatially resolved particle energy distributions $f_{s}(E,x,y)$, we analyze how nonthermal features correlate with shocks and current sheets. On the bottom left side of Fig. 7 we show the local structure of the nonthermal electron $n_{\rm e,nt}$ and nonthermal ion $n_{\rm i,nt}$ particle density. Nonthermal particles are defined here as those with $E>4\langle E_{s}\rangle$, where $\langle E_{s}\rangle$ is the box-averaged species mean energy ($\approx 0.13\,m_{\rm i}c^{2}$ for ions and $\approx 0.19\,m_{\rm e}c^{2}$ for electrons). By comparison with snapshots of $J_{z}$ and of $\nabla\cdot{\boldsymbol{\beta}}_{\rm bulk}$, we find that nonthermal electron populations are enhanced near regions of intense current, whereas the ions display higher nonthermal particle fractions in the vicinity of both shocks and current sheets.
On the bottom right side of Fig. 7 we show examples of local particle distributions in the vicinity of a shock and a current sheet. The local distributions are normalized such that the integral $\int f_{s}(E){\rm d}E$ matches the box-averaged result. The analyzed locations are indicated with red squares (for the current sheet) and green squares (for the shock) in the left panels of Fig. 7. We made sure that the selected point near a shock does not accidentally coincide with a current sheet (and vice versa). The local analysis shows that the electron spectrum near a current sheet can be quite hard (for the particular case shown $p\approx 1.8$). On the other hand, at the location of the shock electrons show no obvious signs of enhanced energization. The ion spectrum shows an above-average amount of energetic particles at both the location of the shock and the current sheet. Particularly at the shock, the local nonthermal ion spectrum is notably harder than the boxed-averaged result. We inspected various locations near shocks and reconnecting current sheets and found local particle distributions broadly similar to those shown on the bottom right side of Fig. 7.
Finally, on the top right side of Fig. 7 we show the energy dependence of the mean cosine of the particle pitch angle $\mu=\langle|p_{\parallel}/p|\rangle$, where $p_{\parallel}$ is the momentum parallel to the local $\mathbf{B}$ field and $p$ is the total momentum. Following previous works (Comisso & Sironi, 2019), we compute $\mu$ in the local $E\times B$ frame. In the nonthermal parts of the distribution, the electrons closely align their momenta with the field-parallel direction. This provides additional evidence that electrons are primarily energized at reconnecting current sheets, and in particular via the nonideal $E_{\parallel}$ fields (see also Comisso & Sironi, 2019; Comisso, 2024). In contrast, the ion momentum distribution is more isotropic and shifts slightly toward lower values of $\mu$ with growing energy.
Appendix B Photon-photon and Ion-electron Collision Rates
Apart from Compton scattering, our simulations do not account for other binary particle interactions, such as Coulomb collisions or two-photon pair production, which may be relevant in black-hole coronae (Fabian et al., 2015; Bambic et al., 2024; Nättilä, 2024). Here we estimate the expected rates of these interactions.
Pair production
With growing compactness $\ell$, the system can transition into a regime with copious electron-positron pair production, which would invalidate our assumed electron-ion composition. The mean steady-state positron density $n_{+}$ can be estimated by balancing pair production and annihilation (Svensson, 1987; Beloborodov, 2017):
$\dot{n}_{\gamma\gamma}\simeq\eta\sigma_{\rm T}cn_{1}^{2}\simeq\dot{n}_{\rm ann}\simeq(3/8)\sigma_{\rm T}cn_{+}(n_{0}+n_{+}),$ (B1)
where $\eta\simeq 0.1$ and $n_{1}=f_{1}n_{\rm ph}$ is the density of photons with energies $m_{\rm e}c^{2}$. By introducing the pair-loading factor $Z_{\pm}=2n_{+}/n_{0}$, we can estimate for $\eta\simeq 0.1$ that
$Z_{\pm}\simeq\left[1+(f_{1}n_{\rm ph}/n_{0})^{2}\right]^{1/2}-1.$ (B2)
Using (12) we find that $Z_{\pm}\gtrsim 1$ when:
$\sigma_{\rm i}\gtrsim\left[\frac{\sqrt{12}E_{0}(A-1)}{f_{1}m_{\rm i}c^{2}(1-q_{\rm i})(\tau_{\rm T}+1)}\right]^{2/3}.$ (B3)
The measured fraction of photons with $E>m_{\rm e}c^{2}$ in our runs with $\ell\approx 0.6,2.5,8.5$ is $f_{1}\approx 3.4\!\times\!10^{-6},6.5\!\times\!10^{-6},7.4\!\times\!10^{-6}$, respectively. With a conservative choice of $f_{1}=10^{-5}$, we find for a representative set of parameters ($m_{\rm i}/m_{\rm e}=1836$, $q_{\rm i}=0.65$, $\tau_{\rm T}=1$, $E_{0}=2.7\times 10^{-5}m_{\rm e}c^{2}$, and A1=16A-1=16) that condition (B3) amounts to $\sigma_{\rm i}\gtrsim 0.24$. According to (2), significant pair production is therefore expected only at the high end of the plausible $\sigma_{\rm i}$ range, where $\ell\sim 100$. Note that the mapping between $\sigma_{\rm i}$ and $\ell$ is artificially modified in our simulations, owing to the reduced ion-electron mass ratio. For this reason, our largest simulated $\ell\approx 8.5$ is far below the range for pair production, even though the corresponding $\sigma_{\rm i}=0.19$ might be relatively close to the $Z_{\pm}\sim 1$ boundary for a realistic $m_{\rm i}/m_{\rm e}$.
Ion-electron collisions
Rapid Coulomb collisions between ions and electrons can lead to equilibration of their temperatures. It is therefore useful to check when collisions are sufficiently infrequent to allow for the development of a two-temperature corona with $T_{\rm i}\gg T_{\rm e}$. For simplicity, we employ analytical estimates for a Maxwellian plasma with dimensionless species temperatures $\theta_{\rm s}=T_{\rm s}/(m_{\rm s}c^{2})$ (Stepney, 1983; Stepney & Guilbert, 1983), where in practice $\theta_{\rm s}$ is to be understood as an effective kinetic temperature of the (nonthermal) particle distribution. The collisional ion-electron relaxation time scale, on which hot ions lose energy to cool electrons, can be obtained as $t_{\rm ie}\simeq\left|\kappa_{\rm e}n_{0}(T_{\rm i}-T_{\rm e})/({\rm d}U_{\rm e}/{\rm d}t)\right|$ (Spitzer, 1962), where ${\rm d}U_{\rm e}/{\rm d}t$ is the collisional electron energy exchange rate per unit volume, and $\kappa_{\rm e}=1/(\Gamma_{\rm e}-1)$. For non-relativistic or very mildly relativistic particles we can approximate (Stepney, 1983):
$\frac{t_{\rm ie}}{s_{\rm esc}/v_{\rm A}}\simeq\frac{\sqrt{\pi}(m_{\rm i}/m_{\rm e})(v_{\rm A}/c)}{\sqrt{2}\tau_{\rm T}\log\Lambda}\left(\theta_{\rm e}+\theta_{\rm i}\right)^{3/2},$ (B4)
where the Coulomb logarithm $\log\Lambda\simeq 20$. Using $\theta_{\rm i}\simeq\sigma_{\rm i}/\Gamma_{i}$ (per Eq. (7)), we find for $\theta_{\rm e}=0.1$ and $m_{\rm i}/m_{\rm e}=1836$ that the ion-electron collisional equilibration time is longer than the dynamic time ($t_{\rm ie}>s_{\rm esc}/v_{\rm A}$) when $\sigma_{\rm i}\gtrsim 0.04$.
Appendix C Turbulent Mode Decomposition
Further insight into the nature of the turbulent cascade can be obtained by decomposing the turbulent-energy fluctuations among different linear modes supported by the plasma. The identification of modes is based on their polarization properties, which are reasonably well preserved even in strongly turbulent regimes (Grošelj et al., 2019). Here, we consider the partitioning of energy between the MHD Alfvén, fast, and slow mode. Apart from its simplistic appeal, the use of MHD relations is supported by studies showing that the large-scale dynamics of turbulent kinetic plasmas is often remarkably MHD-like (e.g., Verscharen et al., 2017; Meyrand et al., 2019; Squire et al., 2019).
Our 2D simulation domain is limited to wave vectors $\mathbf{k}_{\perp}$ perpendicular to the out-of-plane mean magnetic field $\mathbf{B}_{0}=B_{0}\hat{\mathbf{z}}$. However, an effective field-parallel wave number $k_{\parallel}$ is still retained via the projection of $\mathbf{k}_{\perp}$ onto the local magnetic field (e.g., Li et al., 2016; Grošelj et al., 2017). Under these circumstances, it is most reasonable to consider linear modes in the limit of quasi-perpendicular propagation (see Schekochihin, 2025, for details). In this limit, the decomposition becomes independent of $k_{\parallel}$, as shown below. For full 3D mode decomposition at arbitrary $\mathbf{k}$, see e.g. Cho & Lazarian (2003).
To lowest order in $k_{\parallel}/k_{\perp}\ll 1$, the polarizations are determined as follows. For a given $\mathbf{k}\approx\mathbf{k}_{\perp}$, shear Alfvén modes feature magnetic field and velocity fluctuations aligned with the $\mathbf{B}_{0}\times\mathbf{k}_{\perp}$ direction, while the electric field is aligned with $\mathbf{k}_{\perp}$. The fast mode has velocity vector aligned with $\mathbf{k}_{\perp}$, and its electric field is parallel to $\mathbf{B}_{0}\times\mathbf{k}_{\perp}$. The velocity fluctuation of the slow mode is along $\mathbf{B}_{0}$. Fluctuations of the plasma density $\delta n$ and magnetic field magnitude $\delta B\approx\delta B_{z}$ are contained in both the slow and fast mode. Introducing the variables $\delta\tilde{p}=4\pi\delta p/B_{0}$ and $\delta\tilde{n}=B_{0}\delta n/n_{0}$, where $\delta p$ is the pressure fluctuation, we can decompose $\delta B$ as follows:
$\delta B$ $=\delta B_{\rm slow}+\delta B_{\rm fast},$ (C1)
$\delta B_{\rm slow}$ $=(1+1/\alpha)^{-1}(\delta B-\delta\tilde{n})=-\delta\tilde{p}_{\rm slow},$ (C2)
$\delta B_{\rm fast}$ $=(1+\alpha)^{-1}(\delta B+\alpha\delta\tilde{n})=\delta\tilde{n}_{\rm fast},$ (C3)
where the pressure fluctuation is related via the adiabatic response to the density as $\delta\tilde{p}=\alpha\delta\tilde{n}$, $\alpha=c_{\rm s}^{2}/v_{\rm A}^{2}$ is the squared ratio of the sound speed $c_{\rm s}$ to the Alfvén velocity $v_{\rm A}$, the “slow” subscript denotes the slow mode contribution, and the “fast” subscript indicates the part identified as the fast mode. Eq. (C2) shows that the slow mode fluctuations are essentially pressure-balanced, whereas Eq. (C3) implies conservation of magnetic flux by the fast-mode plasma compressions. The turbulent energy at a given $\mathbf{k}_{\perp}$ can be then written as a sum of the 3 modes:
$\mathcal{E}(\mathbf{k}_{\perp})=\,$ $\mathcal{E}_{\rm Alfven}(\mathbf{k}_{\perp})+\mathcal{E}_{\rm slow}(\mathbf{k}_{\perp})+\mathcal{E}_{\rm fast}(\mathbf{k}_{\perp})$
$=\,$ $\frac{1}{8\pi}\left\{|\mathbf{B}(\mathbf{k}_{\perp})|^{2}+|\mathbf{E}_{\perp}(\mathbf{k}_{\perp})|^{2}\right\}$
$+|{\boldsymbol{\upsilon}}(\mathbf{k}_{\perp})|^{2}+\frac{\alpha}{8\pi}|\delta\tilde{n}(\mathbf{k}_{\perp})|^{2},$ (C4)
$\mathcal{E}_{\rm Alfven}(\mathbf{k}_{\perp})=\,$ $\frac{1}{8\pi}|\mathbf{B}_{\perp}(\mathbf{k}_{\perp})|^{2}+|(\hat{\mathbf{z}}\times\hat{\mathbf{k}}_{\perp})\cdot{\boldsymbol{\upsilon}}_{\perp}(\mathbf{k}_{\perp})|^{2}$
$+\frac{1}{8\pi}|\hat{\mathbf{k}}_{\perp}\cdot\mathbf{E}_{\perp}(\mathbf{k}_{\perp})|^{2},$ (C5)
$\mathcal{E}_{\rm slow}(\mathbf{k}_{\perp})=\,$ $\frac{1+1/\alpha}{8\pi}|\delta B_{\rm slow}(\mathbf{k}_{\perp})|^{2}+|\upsilon_{z}(\mathbf{k}_{\perp})|^{2},$ (C6)
$\mathcal{E}_{\rm fast}(\mathbf{k}_{\perp})=\,$ $\frac{1+\alpha}{8\pi}|\delta B_{\rm fast}(\mathbf{k}_{\perp})|^{2}+|\hat{\mathbf{k}}_{\perp}\cdot{\boldsymbol{\upsilon}}_{\perp}(\mathbf{k}_{\perp})|^{2}$
$+|(\hat{\mathbf{z}}\times\hat{\mathbf{k}}_{\perp})\cdot\mathbf{E}_{\perp}(\mathbf{k}_{\perp})|^{2},$ (C7)
where $\hat{\mathbf{k}}_{\perp}=\mathbf{k}_{\perp}/k_{\perp}$, $\boldsymbol{\upsilon}=[\Gamma_{\rm bulk}^{3}w_{0}/(\Gamma_{\rm bulk}+1)]^{1/2}{\boldsymbol{\beta}}_{\rm bulk}$ is the weighted fluid velocity, $\Gamma_{\rm bulk}=[1-\beta_{\rm bulk}^{2}]^{-1/2}$ is the fluid Lorentz factor, and $w_{0}$ is the mean proper enthalpy density. [^8] For completeness, we retain the electric field contribution and we use a relativistic definition of the kinetic energy density, but we note that relativistic effects are mild in the regime of interest. The last term in Eq. (C4) can be written in real space as $\delta n\delta p/(2n_{0})$. It can be identified as the energy density of acoustic oscillations, which naturally enter the total energy definition when the cascade is compressible (e.g., see Schekochihin et al., 2009).
The above-described MHD mode decomposition is applied to our simulation data in spectral space and in real space. Since the ions are much hotter than electrons, the ion pressure and density are used to obtain $\delta\tilde{p}$ and $\delta\tilde{n}$, respectively. The ratio $\alpha$ is estimated at each given time from the root-mean-square fluctuations of the pressure and density as $\alpha=\delta\tilde{p}_{\rm rms}/\delta\tilde{n}_{\rm rms}$. Given the large-amplitude nature of our turbulence drive, the mode decomposition is not exact, even within the scope of MHD. In particular, the direction parallel to the local magnetic field $\hat{\mathbf{b}}=\mathbf{B}/B$ can significantly deviate from $\mathbf{B}_{0}$. In real space, the fluctuating fields can be readily projected onto the local field-parallel direction. Therefore, in real space we use $\delta B\approx\delta(\hat{\mathbf{b}}\cdot\mathbf{B})=\delta|B|$ when extracting the fast mode fluctuations. On the other hand, in spectral space the local projection introduces a certain ambiguity, because it convolves modes with different field components and wavenumbers. We therefore prefer to use $\mathbf{B}_{0}$ as the field-parallel direction for the spectral decomposition. Nevertheless, we confirmed that similar results are obtained when the parallel direction is defined with respect to $\hat{\mathbf{b}}$.